The molar volume (cm3/mol) of a binary liquid mixture at a (40) specific T and P is given by: V = 200x1 + 300x2 + x1 x2 (20x1 + 10x2) (a) Find expressions for the partial molar volumes of species 1 and species 2 at specific T and P (b) Show that the expressions for partial molar volumes follow the summability relation (c) Show that these expressions satisfy the Gibbs/Duhem equation at constant T and P (d) Find the values for the followings – ̅∞ ̅∞ ?;?;?;?

ONLY part d

v1, v2, v1 bar infinity and v2 bar infinity
Answer #1

Sel V = 2ook + 300 ks + koks (don floxo) V=Joox +30o-300x + (4-2) (gox tlo-lon) V= D 300- looni + (x-xi) ( 10 +104) V = 300 -Is Vi - If ru V = lt V = 210-30 + 50 =1200 cm / m Vy = 300+gol 1) = 1390 cm med V = lt Plocm mmal 300 cm / med ابر 110-0+0= J

using given data and required formulae for partial molar properties we can find the pure molar volume and molar volume at infinite dilution

for any doubts write in comments